Session: Feb 1998 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Thu Feb 12 12:22:37 MET 1998) Minor failures
Strong EMI just before 10UT

Medicina (M. Bondi at Thu Feb 12 18:14:36 MET 1998) Minor failures
Problems with tracking the source in a couple of scans

Onsala (John Conway at Fri Feb 13 10:01:22 MET 1998) Minor failures
OK, but we had high Tsys, SEFD 1000 - 1500Jy instead of the the expected 800Jy, we had very heavy rainfall and are investigating whether any water has get into the feed.

Jodrell (Alastair Gunn at Fri Feb 13 11:20:01 MET 1998) Success
No known problems. Clock offset ~7x10-7.

Cambridge (Alastair Gunn at Fri Feb 13 11:21:11 MET 1998) Minor failures
Missed passes 2 and 3 due to vacuum problems. FS terminated and restarted at 06:23. Data should be good from pass 4. Tape started approx. 30s late on pass 24 (tape 1) due to loss of vacuum. At 08:00 BBC04 changed to IF input A from C due to LO signal appearing in band. Clock offset ~4x10^-7.

Torun (K.Borkowski at Fri Feb 13 13:53:24 MET 1998) Success
No known problems.

Westerbork (tony foley at Sat Feb 14 15:48:31 MET 1998) Minor failures
Jump in 1pps at formatter (gps-fmout) from ~ 1.5 to ~2.5 microseconds about 11:10UT. There was no corresponding jump in the maser offset so this must be a formatter jump! I resynched the formatter after the experiment

Effelsberg (dgraham at Mon Feb 16 16:00:52 MET 1998) Minor failures
sources too low until about 0650UT, behind hills.

MPIfR (dgraham at Mon Feb 16 16:13:59 MET 1998) Minor failures
fringes found eff-ons-jod-cbg-wsrt7 on scan 064250 (2134+00), quality OK.

Urumqi (Liu Xiang at Thu Feb 19 04:06:47 MET 1998) Success
It was clear but there was some snow on the surface of the antenna, the clock offset 4.2us+/-0.5us.

Noto (C.Trigilio at Fri Feb 20 15:35:10 MET 1998) Minor failures
Schedule started at 07:17 for an operator problem. No other problems

Calibration (Jean\-Francois Desmurs at Mon Mar 2 18:42:21 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name em25a.antab. Read the comments in the file to get further information.

MPIfR (Walter ALef at Wed May 27 17:23:52 METDST 1998) Minor failures
No fringes for Urumqi and Torun with unkown reason. Noto fringes very weak; looks like wrong polarisation. Cambridge no phase-cal. Crimea missing for the first 4 hours dur to PI error in schedule (procedure made with old version of PC-SCHED). Rest O.K.